OpenStudy (anonymous):

use substitution to show the limit=1 limit [(1/h)/(1+ 1/h)] h tends to zero

OpenStudy (anonymous):

Let \( x=\frac 1 h\) \[ \lim_{h\to 0} (1/h)/(1+ 1/h) =\lim_{x\to \infty} \frac x{1 +x}=1 \]

OpenStudy (anonymous):

Got it thanks

OpenStudy (anonymous):

yw